# Binary Hypothesis Testing-Based Low-Complexity Beamspace Channel Estimation for mmWave Massive MIMO Systems

### v1

@@ Abstract
Index Terms:
Beamspace channel estimation, mmWave, massive MIMO, Bernoulli-complex Gaussian prior, binary hypothesis testing, low-complexity.
@@ I-A Prior Works
Previous works have proposed various channel estimation methods for mmWave MIMO systems. Specifically, several works have utilized methods based on compressive sensing (CS) frameworks that exploit the beamspace sparsity of mmWave channels[14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The works in [14, 15, 16] proposed channel estimation frameworks using approximate message passing (AMP), which is a representative CS algorithm. AMP-based channel estimation algorithms have also been developed for low-resolution receivers[17, 18]. Meanwhile, orthogonal matching pursuit (OMP), another representative CS algorithm, has also been adopted for channel estimation[19, 20, 21, 22, 23]. However, these algorithms require iterative optimization or are computational intensive, which makes their adoption difficult in dynamic propagation environments that change rapidly.
@@
In addition, several previous works applied deep learning for mmWave channel estimation[24, 25, 26, 27, 28, 29]. The method in [24] proposed a supervised denoising model based on AMP and a denoising convolutional neural network (DnCNN) for beamspace channel estimation. Ma et al. devised a learned AMP-based sparse channel estimation framework, which reduces pilot overhead by jointly optimizing the pilot sequence and the channel estimator[25]. Abdallah et al. proposed a DnCNN-based channel estimator that learns channel support and amplitudes in the frequency domain[26]. Wu et al. applied deep unfolding for sparse beamspace channel estimation leveraging trimmed-ridge regression[27]. Jang and Lee proposed the joint optimization of environment-specific beamformers and geometric parameters based on a neural network for wideband mmWave channel estimation[28]. Zhang et al. proposed a denoising approach using a generative adversarial network (GAN), which refines noisy channel estimates by leveraging spatial and temporal correlations in time-varying mmWave channels[29]. However, aforementioned frameworks require paired ground-truth data, implying a strong dependence on potentially expensive datasets containing a large number of noiseless channel realizations for training. To overcome this limitation, unsupervised learning-based channel estimation methods have been devised[30, 31, 32, 34, 33]. Tong et al. proposed an unsupervised learning-based two-step OMP algorithm for channel estimation[30]. The work in [31] adopted subcarrier-wise estimation of signal parameters via rotational invariant techniques (ESPRIT) and $k$-means clustering to estimate the wideband mmWave channel. Gahlot et al. devised a training-free channel estimation that performs denoising using anisotropic diffusion model[32]. He et al. formulated wideband beamspace channel estimation as a CS problem on the 2-dimensional (2D) image using deep unfolding, and applied Stein’s unbiased risk estimate (SURE) to enable unsupervised learning[33]. In [34], the authors proposed deep equilibrium model (DEQ) for beamspace channel estimation, which eliminates the dependency on ground-truth data and the white noise assumption inherent in SURE by using generalized SURE (GSURE). Nevertheless, these methods require training phase and have weak explainability. Moreover, the substantial computational demands of deep learing-based methods may hinder their practicality in realistic wireless environments and hardware-constrained deployments.
@@
To reduce computational burden, several studies have proposed frameworks that reduce complexity[35, 36, 37, 38, 39, 40, 41, 42]. The works in [35] and [36] proposed frameworks that reduce the complexity of sparse Bayesian learning-based channel estimation. Zhou et al. utilized matrix factorization and the log-barrier function to reduce the complexity of sparse channel estimation for one-bit receivers[37]. Masood et al. applied low-rank matrix sensing problem and low-rank matrix completion for channel estimation, exploiting channel sparsity to reduce the computational complexity[38]. Ruan et al. proposed a GAN-based beamspace channel estimation algorithm that reduces training complexity by leveraging Wasserstein GAN with simplified learned AMP (LAMP)[39]. The work in [40] employed subspace-based channel estimation and simplified the process using a low-dimensional least squares problem. Zhu et al. applied semidefinite programming and alternating direction method of multipliers (ADMM) to simplify the original atomic norm minimization problem of channel estimation, leveraging Hankel-Toeplitz structures[41]. In [42], a soft-thresholding-based beamspace channel estimation algorithm was devised, which adaptively denoises the noisy channel using SURE with a limited number of iterations. However, these algorithms still involve iterative search, optimization, or learning phases, leaving room for further improvement in simplicity of implementation and computational efficiency.
@@ I-B Contributions
Since mmWave channels exhibit rapid fluctuations, fast and computationally efficient channel estimation becomes essential. However, the complexity of existing estimation algorithms typically scales with the number of antennas and further increases with iterative search or learning, which poses significant challenges in terms of latency and computing resource requirements in large-scale mmWave MIMO systems. To address this challenge, in this paper, we propose a novel beamspace channel denoiser for mmWave massive MIMO systems, which does not rely on iterative optimization and learning-based approach. It achieves low computational complexity by leveraging Bayesian binary hypothesis testing under Bernoulli-complex Gaussian distribution, which effectively captures the sparsity of beamspace channel model. The detailed contributions of this work are summarized as follows:
@@ I-C Paper Outline
[…]
@@
Also, $\mathds{1}_{\{\mathcal{A}\}}$ represents the indicator function with event $\mathcal{A}$, which is defined as
$\mathds{1}_{\{\mathcal{A}\}}=\begin{cases}1&\textnormal{if }\mathcal{A}\textnormal{ is true,}\\ 0&\textnormal{otherwise.}\end{cases}$ (2)
@@ II System Model
In the following, we remove the user index for notational convenience. Without loss of generality, we assume that the channel is observed with orthonormal pilot sequences, and the BS estimates the channel vector for each user. With uplink pilot, the BS can obtain a noisy channel observation via least-squares (LS) estimation, which is presented as
@@
where $\mathbf{e}\sim\mathcal{CN}(0,E_{0}\mathbf{I}_{M})$ is the channel estimation error. Since we assumed that the pilot is orthonormal, $E_{0}$ is equivalent to $N_{0}$. To perform beamspace processing, the observed channel is transformed from the antenna domain into beamspace, which can be expressed as
$\bar{\mathbf{h}}^{\prime}=\mathbf{Fh}+\mathbf{Fe}=\bar{\mathbf{h}}+\bar{\mathbf{e}},$ (7)
where $\bar{\mathbf{h}}$ and $\bar{\mathbf{e}}$ are the ground-truth channel and channel estimation error in beamspace domain, respectively, and $\mathbf{F}\in\mathbb{C}^{M\times M}$ is the normalized discrete Fourier transform matrix implying $\mathbf{F}^{H}=\mathbf{F}^{-1}$. Note that $\bar{\mathbf{e}}$ is still AWGN, and $\|\mathbf{h}\|=\|\bar{\mathbf{h}}\|$. Due to the limited scattering in mmWave propagation environments, noiseless channel vector $\bar{\mathbf{h}}$ is approximately sparse. This assumption can be reformulated using Bernoulli-complex Gaussian distribution[17, 18, 39, 44], which implies that the PDF of noisy beamspace channel vector element is determined as
$f_{\bar{\mathbf{h}}}(\bar{h}_{m})\triangleq$ $q\frac{1}{\pi(E_{0}+\|\bar{\mathbf{h}}\|^{2}/q)}\exp\left(-\frac{|\bar{h}_{m}|^{2}}{E_{0}+\|\bar{\mathbf{h}}\|^{2}/q}\right)$ (8)
$+(1-q)\delta(\bar{h}_{m}),$
where $q=\|\bar{\mathbf{h}}\|_{0}/M$ is activity rate of the channel and $\delta(\cdot)$ is the Dirac-delta distribution. To provide intuition, the activity rate $q$ indicates the fraction of active beams in the beamspace channel vector. For example, in a massive MIMO system with $M=128$ antennas, if only 20 elements in the noiseless channel vector carry signal energy and other elements has near-zero magnitudes, then $q=20/128=0.15625$. This reflects the inherent sparsity of mmWave channels in beamspace caused by limited scattering, where only a small subset of angular directions contains significant multipath components. Based on this modeling, the noisy observation of the mmWave beamspace channel can be derived as
@@
[…]
@@ III Hypothesis Testing-Based Beamspace Channel Estimation
Based on this testing, Bayesian hypothesis testing based on likelihood ratio test can be organized as
@@
$|\bar{h}_{m}|^{2}$ $\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}N_{0}\left(\frac{p(\mathcal{H}_{1})}{\mathrm{SNR}}+1\right)$ (15)
×$\quad\quad\times\ln\left(\left(1+\frac{\mathrm{SNR}}{p(\mathcal{H}_{1})}\right)\times\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right),$
where $\mathrm{SNR}=\|\bar{\mathbf{h}}\|^{2}/N_{0}=\|\bar{\mathbf{h}}\|^{2}/E_{0}$. For the complete derivation, please refer to Appendix A. This threshold increases with the noise power and signal sparsity, intuitively suppressing components less likely to be active.
@@
The median can be found at a complexity of $\mathcal{O}(M)$ time either in both the worst and average cases using median of medians[45] or on average using the quickselect algorithm[46]. Also, it is also used in blind SNR estimation, which is given by
@@
$\mu_{4}$ $\triangleq\mathbb{E}[|\bar{\mathbf{h}}^{\prime}|^{4}]$ (18)
$=p(\mathcal{H}_{0})\mathbb{E}[|\bar{\mathbf{h}}^{\prime}|^{4}|\mathcal{H}_{0}]+p(\mathcal{H}_{1})\mathbb{E}[|\bar{\mathbf{h}}^{\prime}|^{4}|\mathcal{H}_{1}],$
with the law of total expectation. First, since $\bar{\mathbf{h}}^{\prime}|\mathcal{H}_{0}\sim\mathcal{CN}(0,E_{0})$, its fourth-order moment is given by
@@
which is derived from the well-known result of fourth-order moment of zero-mean complex Gaussian. Likewise, the fourth-order moment in $\mathcal{H}_{1}$ is determined as
@@
$q$ $=\frac{2\|\mathbf{h}\|^{4}}{\mu_{4}-2E_{0}^{2}-4E_{0}\|\mathbf{h}\|^{2}}$ (22)
$=\dfrac{2(\mathrm{SNR})^{2}}{\dfrac{\mu_{4}}{E_{0}^{2}}-2-4\mathrm{SNR}}.$
@@
and thus
$\mathrm{Var}(\hat{q})\propto q^{2}.$ (34)
@@
In summary, the proposed method performs signal denoising with five main steps: 1) blind noise power estimation, 2) blind SNR estimation, 3) blind activity rate estimation, 4) threshold calculation, and 5) hard thresholding. In terms of computational complexity, blind estimators for noise power, SNR, and blind activity rate estimation require $\mathcal{O}(M)$ for each. Also, since threshold calculation is derived as a closed-form, it only requires $\mathcal{O}(1)$. The final step, hard thresholding, requires $\mathcal{O}(M)$ to scan all elements of beamspace channel vector and perform detection. Overall, the estimator only requires $\mathcal{O}(M)$, which represents extremely reduced complexity, which is discussed in detail compared to other channel estimation algorithms in Section IV.

### v2

@@ I-A Prior Works
Previous works have proposed various channel estimation methods for mmWave MIMO systems. These works can be broadly grouped into three directions. The first line of research exploits beamspace sparsity of mmWave channels through compressive sensing (CS) frameworks [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The works in [14, 15, 16] proposed channel estimation frameworks using approximate message passing (AMP), which is a representative CS algorithm. AMP-based channel estimation algorithms have also been developed for low-resolution receivers[17, 18]. Meanwhile, orthogonal matching pursuit (OMP), another representative CS algorithm, has also been adopted for channel estimation[19, 20, 21, 22, 23]. However, these algorithms require iterative optimization or are computationally intensive, which makes their adoption difficult in dynamic propagation environments that change rapidly.
@@
The second research trend adopts deep learning for channel estimation[24, 25, 26, 27, 28, 29]. The method in [24] proposed a supervised denoising model based on AMP and a denoising convolutional neural network (DnCNN) for beamspace channel estimation. Ma et al. devised a learned AMP-based sparse channel estimation framework, which reduces pilot overhead by jointly optimizing the pilot sequence and the channel estimator[25]. Abdallah et al. proposed a DnCNN-based channel estimator that learns channel support and amplitudes in the frequency domain[26]. Wu et al. applied deep unfolding for sparse beamspace channel estimation leveraging trimmed-ridge regression[27]. Jang and Lee proposed the joint optimization of environment-specific beamformers and geometric parameters based on a neural network for wideband mmWave channel estimation[28]. Zhang et al. proposed a denoising approach using a generative adversarial network (GAN), which refines noisy channel estimates by leveraging spatial and temporal correlations in time-varying mmWave channels[29]. However, aforementioned frameworks require paired ground-truth data, implying a strong dependence on potentially expensive datasets containing a large number of noiseless channel realizations for training. To overcome this limitation, unsupervised learning-based channel estimation methods have been devised[30, 31, 32, 33, 34, 36, 35]. Tong et al. proposed an unsupervised learning-based two-step OMP algorithm for channel estimation[30]. The work in [31] adopted subcarrier-wise estimation of signal parameters via rotational invariant techniques (ESPRIT) and $k$-means clustering to estimate the wideband mmWave channel. Gahlot et al. devised a training-free channel estimation that performs denoising using anisotropic diffusion model[32]. Balevi et al. proposed a CS-based channel estimation with deep generative network[33]. Zhou et al. utilized a diffusion model in MIMO channel estimation problem with Stein’s unbiased risk estimate (SURE)[34]. He et al. formulated wideband beamspace channel estimation as a CS problem on the 2-dimensional (2D) image using deep unfolding, and applied SURE to enable unsupervised learning[35]. In [36], the authors proposed deep equilibrium model (DEQ) for beamspace channel estimation, which eliminates the dependency on ground-truth data and the white noise assumption inherent in SURE by using generalized SURE (GSURE). Nevertheless, these methods require training phase and have weak explainability. Moreover, the substantial computational demands of deep learing-based methods may hinder their practicality in realistic wireless environments and hardware-constrained deployments.
@@
Another line of research focuses on reducing computational complexity through lightweight denoisers and simplified frameworks, as the aforementioned methods can become computationally demanding when the number of antennas increases[37, 38, 39, 40, 41, 42, 43, 44, 45]. The works in [37] and [38] proposed frameworks that reduce the complexity of sparse Bayesian learning-based channel estimation. Zhou et al. utilized matrix factorization and the log-barrier function to reduce the complexity of sparse channel estimation for one-bit receivers[39]. Masood et al. applied low-rank matrix sensing problem and low-rank matrix completion for channel estimation, exploiting channel sparsity to reduce the computational complexity[40]. Ruan et al. proposed a GAN-based beamspace channel estimation algorithm that reduces training complexity by leveraging Wasserstein GAN with simplified learned AMP (LAMP)[41]. The work in [42] employed subspace-based channel estimation and simplified the process using a low-dimensional least squares problem. Zhu et al. applied semidefinite programming and alternating direction method of multipliers (ADMM) to simplify the original atomic norm minimization problem of channel estimation, leveraging Hankel-Toeplitz structures[43]. Fesl et al. devised a lightweight CNN within a diffusion-based generative prior to achieve low-complexity MIMO channel estimation[44]. In [45], a soft-thresholding-based beamspace channel estimation algorithm was devised, which adaptively denoises the noisy channel using SURE with a limited number of iterations. However, these algorithms still involve iterative search, optimization, or learning phases, leaving room for further improvement in simplicity of implementation and computational efficiency.
@@ I-B Contributions
Since mmWave channels exhibit rapid fluctuations[7], fast and computationally efficient channel estimation becomes essential. However, the complexity of existing estimation algorithms typically scales with the number of antennas and further increases with iterative search or learning, which poses significant challenges in terms of latency and computing resource requirements in large-scale mmWave MIMO systems. To address this challenge, in this paper, we propose a novel beamspace channel denoiser for mmWave massive MIMO systems, which does not rely on iterative optimization and a learning-based approach. It achieves low computational complexity by leveraging Bayesian binary hypothesis testing under Bernoulli-complex Gaussian distribution, which effectively captures the sparsity of beamspace channel model. The detailed contributions of this work are summarized as follows:
@@ I-C Paper Outline
Matrices and vectors are denoted by boldface uppercase and lowercase letters, respectively. All vectors are column vectors. For matrix $\mathbf{A}$, its $i$th column vector is $\mathbf{a}_{i}$, and its $j$th element is $A_{ij}$. $(\cdot)^{T}$ means transpose, $(\cdot)^{H}$ denotes conjugate transpose, and $\mathbf{I}_{N}$ is $N\times N$ identity matrix. $\mathbf{0}^{N\times M}$ is $N\times M$ zero matrix. $\|\cdot\|$ represents the $\ell_{2}$ norm, and $\|\mathbf{a}\|_{0}$ is $\ell_{0}$ norms, i.e., the number of nonzero entries in $\mathbf{a}$. Also, $|\mathbf{a}|$ means the element-wise absolute-valued vector for $\mathbf{a}$. $\mathrm{Exp}(\lambda)$ denotes exponential distribution with rate parameter $\lambda$, with the probability density function (PDF)
@@
$\overset{D}{\longrightarrow}$ and $\overset{\textnormal{a.s.}}{\longrightarrow}$ denote convergence in distribution and convergence almost surely, respectively. Also, $\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}$ denotes a binary decision operator in which the decision is $\mathcal{H}_{1}$ if the left-hand side is greater than the right-hand side, and the decision is $\mathcal{H}_{0}$ otherwise.
@@ II System Model
In the following, we remove the user index for notational convenience. Without loss of generality, we assume that the variance of pilot and data symbols, which is denoted by $P_{x}$, is 1. With uplink pilot, the BS can obtain a noisy channel observation via dedicated training phase for each UE, which is presented as
@@
where $\mathbf{e}\sim\mathcal{CN}(0,E_{0}\mathbf{I}_{M})$ is the channel estimation error. Since we assumed that the variance of the pilot is 1, $E_{0}$ is equivalent to $N_{0}$. To perform beamspace processing, the observed channel is transformed from the antenna domain into beamspace, which can be expressed as
$\bar{\mathbf{h}}^{\prime}=\mathbf{Fh}^{\prime}=\mathbf{Fh}+\mathbf{Fe}=\bar{\mathbf{h}}+\bar{\mathbf{e}},$ (6)
where $\bar{\mathbf{h}}$ and $\bar{\mathbf{e}}$ are the ground-truth channel and channel estimation error in beamspace domain, respectively, and $\mathbf{F}\in\mathbb{C}^{M\times M}$ is the normalized discrete Fourier transform matrix implying $\mathbf{F}^{H}=\mathbf{F}^{-1}$. Note that $\bar{\mathbf{e}}$ is still AWGN, and $\|\mathbf{h}\|=\|\bar{\mathbf{h}}\|$. Due to the limited scattering in mmWave propagation environments, noiseless channel vector $\bar{\mathbf{h}}$ is approximately sparse. Using this assumption, the sparsity of the beamspace channel can be simpilified with Bernoulli-complex Gaussian distribution[17, 18, 41, 47], which implies that the PDF of noiseless beamspace channel vector element is determined as
$f_{\bar{\mathbf{h}}}(\bar{h}_{m})\triangleq~q\frac{1}{\pi\|\bar{\mathbf{h}}\|^{2}/q}\exp\left(-\frac{|\bar{h}_{m}|^{2}}{\|\bar{\mathbf{h}}\|^{2}/q}\right)+(1-q)\delta(\bar{h}_{m}),$ (7)
where $q=\|\bar{\mathbf{h}}\|_{0}/M$ is activity rate of the channel and $\delta(\cdot)$ is the Dirac-delta distribution. To provide intuition, the activity rate $q$ indicates the fraction of active beams in the beamspace channel vector. For example, in a massive MIMO system with $M=128$ antennas, if only 20 elements in the noiseless beamspace channel vector $\bar{\mathbf{h}}$ carry signal energy and other elements has near-zero magnitudes, then $q=20/128=0.15625$. This reflects the inherent sparsity of mmWave channels in beamspace caused by limited scattering, where only a small subset of angular directions contains significant multipath components. Based on this modeling, the noisy observation of the mmWave beamspace channel can be derived as
@@
The PDF in (8) models the noisy observation by assigning each beam a probability $q$ of being signal-bearing and (1q)(1-q) of being noise-only.
@@ III Hypothesis Testing-Based Beamspace Channel Estimation
Based on this testing, Bayesian hypothesis testing can be organized as
@@
$|\bar{h}_{m}^{\prime}|^{2}\underset{\mathcal{H}_{0}}{\overset{\mathcal{H}_{1}}{\gtrless}}E_{0}\left(\frac{p(\mathcal{H}_{1})}{\mathrm{SNR}}+1\right)\ln\left(\left(1+\frac{\mathrm{SNR}}{p(\mathcal{H}_{1})}\right)\frac{1-p(\mathcal{H}_{1})}{p(\mathcal{H}_{1})}C\right),$ (14)
where $\mathrm{SNR}=\|\bar{\mathbf{h}}\|^{2}/E_{0}=\|\mathbf{h}\|^{2}/E_{0}$. For the complete derivation, please refer to Appendix A. This threshold increases with the noise power and signal sparsity, intuitively suppressing components less likely to be active.
@@
The median can be found at a complexity of $\mathcal{O}(M)$ time either in both the worst and average cases using median of medians[48] or on average using the quickselect algorithm[49]. Also, it is used in blind SNR estimation, which is given by
@@
$\mu_{4}\triangleq\mathbb{E}[|\bar{\mathbf{h}}^{\prime}|^{4}]=p(\mathcal{H}_{0})\mathbb{E}[|\bar{\mathbf{h}}^{\prime}|^{4}|\mathcal{H}_{0}]+p(\mathcal{H}_{1})\mathbb{E}[|\bar{\mathbf{h}}^{\prime}|^{4}|\mathcal{H}_{1}],$ (17)
with the law of total expectation. First, in Bernoulli-complex Gaussian model, elements of $\bar{\mathbf{h}}$ corresponding to $\mathcal{H}_{0}$ are 0, so they only contain noise in the noisy observation $\bar{\mathbf{h}}^{\prime}$. Thus, $\bar{h}_{m}^{\prime}|\mathcal{H}_{0}\sim\mathcal{CN}(0,E_{0})$, and its fourth-order moment is given by
@@
which is derived from the well-known result of fourth-order moment of zero-mean complex Gaussian. Likewise, the elements of $\bar{\mathbf{h}}$ corresponding to $\mathcal{H}_{1}$ follow Gaussian distribution, its noisy observation is also Gaussian, i.e., $\bar{h}_{m}^{\prime}|\mathcal{H}_{1}\sim\mathcal{CN}(0,E_{0}+\|\bar{\mathbf{h}}\|^{2}/q)$. Thus, the fourth-order moment in $\mathcal{H}_{1}$ is determined as
@@
$q=\frac{2\|\mathbf{h}\|^{4}}{\mu_{4}-2E_{0}^{2}-4E_{0}\|\mathbf{h}\|^{2}}=\dfrac{2(\mathrm{SNR})^{2}}{\dfrac{\mu_{4}}{E_{0}^{2}}-2-4\mathrm{SNR}}.$ (21)
@@
and thus $\mathrm{Var}(\hat{q})\propto q^{2}.$
@@
In summary, the proposed method performs signal denoising with five main steps, which is shown in Algorithm 1: 1) blind noise power estimation, 2) blind SNR estimation, 3) blind activity rate estimation, 4) threshold calculation, and 5) hard thresholding. In terms of computational complexity, blind estimators for noise power, SNR, and blind activity rate estimation require $\mathcal{O}(M)$ for each. Also, since threshold calculation is derived as a closed-form, it only requires $\mathcal{O}(1)$. The final step, hard thresholding, requires $\mathcal{O}(M)$ to scan all elements of beamspace channel vector and perform detection. Overall, the estimator only requires $\mathcal{O}(M)$, which represents extremely reduced complexity, which is discussed in detail compared to other channel estimation algorithms in Section IV.
1input: Noisy beamspace channel vector $\bar{\mathbf{h}}^{\prime}$, testing parameter $C$
2 initialization: estimated beamspace channel vector $\bar{\mathbf{h}}^{\star}=\mathbf{0}^{M\times 1}$
3 Calculate the noise power estimate $\widehat{E}_{0}$
4 Compute $\widehat{\mathrm{SNR}}$ and $\hat{q}$ using $\widehat{E}_{0}$
5 Obtain the threshold with $\widehat{E}_{0},\widehat{\mathrm{SNR}},\hat{q}$ as
$\tau=\widehat{E}_{0}\left(\frac{\hat{q}}{\widehat{\mathrm{SNR}}}+1\right)\ln\left(\left(1+\frac{\widehat{\mathrm{SNR}}}{\hat{q}}\right)\frac{1-\hat{q}}{\hat{q}}C\right)$ (33)
6 for $m=1:M$ do
7 if $|\bar{h}^{\prime}_{m}|^{2}\geq\tau$ then
8    $\bar{h}_{m}^{\star}=\bar{h}_{m}^{\prime}$.
9   end if
10
11 end for
return beamspace channel estimate $\bar{\mathbf{h}}^{\star}$
Algorithm 1 Hypothesis Testing-Based Beamspace Channel Estimation